Fermi National Laboratory


Accelerator Update

Friday March 26
The day shift began with the Antiproton Source (Pbar) stacking, with Main Injector (MI), Recycler, and MiniBooNE experts conducting studies, with the Collider Detector at Fermilab (CDF) and D-zero (D0) in access, and with Switchyard (SY) off due to a power supply problem.

At 9:57 AM, the D0 experimenters reported that the Alignment crew had completed their collision hall alignment work.

D0 and CDF were out of access by at 10:30 AM.

A number of vacuum pumps tripped off in the A1 transfer line at 10:36 AM. Everything reset okay.

Operator discovered a number of MI vacuum valves closed at 10:57 AM. After some investigation operators discovered that the valves had closed because a SY Septa had failed to move to its required position.

At 12:54 PM, a MI expert worked on a MI power supply (QXR), which resulted in beam making it to the SY dump.

At 2:21 PM, an Accumulator power supply (A:LQ) tripped off and terminated the Pbar stack. A technician investigated and discovered loose connections on the supply's breaker. He had the supply repaired by 3:50 PM.

Pbar resumed stacking at 4:34 PM.

Saturday March 27
The midnight shift began with Pbar stacking, with MiniBooNE taking beam, and with Tevatron (TeV) and Recycler experts conducting studies.

At 1:55 AM, Recycler experts concluded their studies.

Operations began experiencing problems with the Tevatron Electron Lens (TEL) at 4:30 AM. Operators called in an expert and with his help managed to correct a timing problem. Operations put the TeV through a dry squeeze at 7:28 AM.

A Booster RF station (BRF16) tripped off at 7:28 AM. It reset okay, but then tripped off ten minutes later. It again tripped off at 7:42 AM and wouldn't reset. Operators bypassed the station.

Operators prepared SY enclosures for access at 7:42 AM for work.

Pbar stopped stacking at 8:39 AM so that TeV experts could conduct TEL studies.

At 10:44 AM, the TeV suffered a sector A1 quench during a manual abort.

At 5:20 PM, TeV studiers tried to load a 36x0 study-store, but couldn't. After much investigation it was determined to be a source code problem.

The TeV suffered a sector F3 quench at 6:45 PM.

At 6:31 PM, operators called in a MI expert to help with a coalescing problem.

Pbar resumed stacking at 6:49 PM.

At 7:20 PM, the MI expert said she'd found a problem with the dampers. Operators called in a damper expert. He discovered the dampers in a wrong state, fixed them, and had coalescing looking better by 7:51 PM.

At 8:30 PM, the TeV suffered a quench due to two bunches being loaded on top of each other.

Operations put the TeV through a dry squeeze at 9:08 PM.

Operations established a 36x0 study-store at 10:19 PM.

Operations put the TeV into shot setup at 11:14 PM.

Sunday March 28
The midnight shift began with the TeV in shot setup, with beam to SY Meson Center, and with beam to MiniBooNE.

At 1:21 AM, Operations reported that the first transfer of antiprotons didn't make it into the TeV. Operators determined that this was a beam-sync problem.

Operations established store 3326 at 2:02 AM with an initial luminosity of 42.96E30. Pbar resumed stacking.

At 9:01 AM, the MCenter experiment stopped wanting beam and so SY beam was terminated. Operations reestablished beam to the SY dump for studies at 11:42 AM.

A MI expert began a four-hour 27 GeV slow acceleration study at 2:45 PM.

At 11:58 PM, a Booster RF station (BRF14) tripped off for the first of three times in two and a half hours.

Monday March 29
The midnight shift began with Operations monitoring store 3326, with Pbar stacking, with SY and Recycler studies, with experts investigating a database problem, and with MiniBooNE taking beam.

At 1:49 AM, Operations terminated store 3326.

Operations put the TeV into shot setup around 2 AM.

Operations established store 3328 at 4:18 AM with an initial luminosity of 42.58E30. Pbar resumed stacking.

Today's Plans
The plan for today is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/29/2004   email Fermilab